Find the x-intercepts

y = x^2 + x - 6
Thank You!

Answer by venugopalramana(3286)

y = x^2 + x - 6 ..TO FIND X INTERCEPTS ,WE HAVE TO FIND WHERE THE CURVE CUTS X AXIS
Y=0 WHEN THE CURVE CUTS X AXIS
X^2+X-6=0=X^2+3X-2X-6=X(X+3)-2(X+3)=0
(X+3)(X-2)=0
X=-3 AND X=2 ARE THE X INTERCEPTS